How do you solve a quadratic equation when there is no b value? e.g. 4x² - 6 = 0
Answer by mananth(16946)

4x² - 6 = 0
add 6
4x^2=6
/4
x^2=6/4 = 3/2
x= +/- sqrt(3/2)